# ReferDINO: Referring Video Object Segmentation with Visual Grounding Foundations

### v1

@@ Abstract
Referring video object segmentation (RVOS) aims to segment target objects throughout a video based on a text description. Despite notable progress in recent years, current RVOS models remain struggle to handle complicated object descriptions due to their limited video-language understanding. To address this limitation, we present ReferDINO, an end-to-end RVOS model that inherits strong vision-language understanding from the pretrained visual grounding foundation models, and is further endowed with effective temporal understanding and object segmentation capabilities. In ReferDINO, we contribute three technical innovations for effectively adapting the foundation models to RVOS: 1) an object-consistent temporal enhancer that capitalizes on the pretrained object-text representations to enhance temporal understanding and object consistency; 2) a grounding-guided deformable mask decoder that integrates text and grounding conditions to generate accurate object masks; 3) a confidence-aware query pruning strategy that significantly improves the object decoding efficiency without compromising performance. We conduct extensive experiments on five public RVOS benchmarks to demonstrate that our proposed ReferDINO outperforms state-of-the-art methods significantly. Project page: https://isee-laboratory.github.io/ReferDINO.
@@ 1 Introduction
Figure 1: (a) The SOTA method [22] fails to handle the descriptions involving composite object attributes and spatial locations. (b) GroundingDINO [21] can neither understand the temporal dynamics nor perform pixel-wise segmentation.
@@
Despite significant progress in recent years, existing RVOS models [2, 35, 22, 24, 37, 8] still suffer from many common problems. For example, they often struggle to handle the descriptions involving complicated object attributes and spatial locations, as shown in Figure 1 (a). These problems are mainly caused by the insufficient vision-language capabilities of current models, which, in turn, result from the limited scale and diversity of available RVOS data.
In this work, we propose ReferDINO, an end-to-end RVOS approach that effectively addresses the above problems by inheriting strong vision-language understanding and spatial grounding capabilities from foundational visual grounding models. Represented by GroundingDINO [21], these foundation models benefit from extensive image-text pretraining and showcase strong potential in understanding complex object-text associations.
The critical challenges in adapting visual grounding models to the RVOS task are the lack of temporal understanding and pixel-wise segmentation capabilities, as illustrated in Figure 1 (b). To bridge these gaps and enable efficient adaptation to RVOS, we propose three key innovations in ReferDINO. First, we introduce an object-consistent temporal enhancer that leverages cross-modal text representations from foundational models to facilitate cross-frame object interaction, improving both temporal understanding and object consistency. Second, we elaborate a grounding-guided deformable mask decoder, which integrates both text and grounding conditions through efficient attention mechanisms to achieve accurate object segmentation. Third, we design a confidence-aware query pruning strategy to alleviate the computational bottlenecks without compromising performance, enabling our ReferDINO to satisfy the efficiency demands of video tasks.
These designs fully unleash the spatial grounding knowledge of pretrained foundation models and endow ReferDINO with effective temporal understanding and object segmentation capabilities. As shown in Figure 1, our ReferDINO overcomes the limitations of state-of-the-art RVOS models and the visual grounding models, enabling accurate referring video object segmentation.
Extensive experiments on five public RVOS benchmarks empirically show remarkable improvements of ReferDINO over state-of-the-art (SOTA) methods. For example, on Ref-DAVIS17 dataset, ReferDINO with the Swin-B backbone outperforms SOTAs by 4.0% $\mathcal{J}\&\mathcal{F}$. Moreover, compared to the competitive baseline [1] that combines GroundingDINO and SAM2 [27] for RVOS, ReferDINO demonstrates significant advantages in performance. For example, on Ref-Youtube-VOS dataset, ReferDINO with the Swin-T backbone improves the baseline by 12.0% $\mathcal{J}\&\mathcal{F}$. These results confirm the effectiveness of our designs over the foundation models. We summarize our contributions as follows:
  • To the best of our knowledge, ReferDINO is the first end-to-end approach for adapting foundational visual grounding models to RVOS.
  • We introduce three key innovations in ReferDINO to enable efficient object decoding, effective temporal understanding and accurate object segmentation.
  • Our ReferDINO significantly outperforms SOTA methods across five public RVOS benchmarks.
@@ 2 Related Works
Figure 2: Overall architecture of ReferDINO. Modules colored in blue are borrowed from GroundingDINO, while those in red are newly introduced in this work. Based on the frame-wise object features $\{\mathcal{O}^{t}\}^{T}_{t=1}$, our object-consistent temporal enhancer leverages the cross-modal text features to enable inter-frame object interaction. Then, our grounding-guided deformable mask decoder produces masks for the candidate objects, conditioned on location predictions, cross-modal text features and high-resolution feature maps. To further improve the efficiency in video processing, we introduce a confidence-aware query pruning strategy in the cross-modal decoder. Best view in color.
@@
Referring Video Object Segmentation. RVOS [6, 5, 30] aims to segment objects throughout the video based on text descriptions. MTTR [2] firstly introduces the DETR paradigm [3] into RVOS. Furthermore, ReferFormer [35] proposes to produce the queries from the text description. On the top of this pipeline, follow-up works [7, 24, 22, 37] focus on modular improvements to improve cross-frame consistency and temporal understanding. Despite notable progress on specific datasets, these models are limited by insufficient vision-language understanding, and often struggle in unseen objects or scenarios. Recently, some works [1, 12] attempt to utilize GroundingDINO to identify objects in single frames and then apply independent segmentation models like SAM2 [27] to generate high-quality object masks. However, such a manner of model ensemble is inefficient and non-differentiable. Meanwhile, their performance almost completely relies on the static detection quality of GroundingDINO. In contrast, our ReferDINO is an end-to-end adaptation approach that benefits from both open-world knowledge of GroundingDINO and specific knowledge from RVOS data.
Visual Grounding Foundation Model. The aim of visual grounding [9, 11, 20, 29, 32, 33] is to locate the most relevant object or region in an image, based on a natural language query. Recent foundational works aim to unify detection datasets and image-text datasets through region-text matching, and exploit large image-text data to increase the training vocabulary at scale. GLIP [19] formulates object detection as a grounding problem, utilizing additional grounding data to promote hierarchical vision-language alignment. GroundingDINO [21] incorpoarates large-scale grounded pretraining into a strong object detector DINO [38] with deep cross-modality fusion. These models enable to capture object locations, attributions and variations based on natural language, and have been preliminary explored in some downstream tasks. For example, Video-GroundingDINO [34] enables GroundingDINO to predict event temporal boundaries by inserting a simple temporal self-attention module. Grounded-SAM [28] directly connects GroundingDINO with SAM [16] to achieve text-prompted image segmentation. In this work, we explore to adapt GroundingDINO for the RVOS task, and propose three critical innovations for efficient adaptation.
@@ 3 Background: GroundingDINO
Our approach builds on a visual grounding foundation model GroundingDINO [21], which we briefly revisit here. GroundingDINO is a DETR-architectural object detector, which introduces language to an object detector to achieve visual grounding. It mainly consists of an image backbone, a text backbone, a cross-modal encoder-decoder Transformer architecture, a box head and a classification head.
Given a pair of image and text, GroundingDINO adopts the dual backbones to extract vanilla features, and then input them into the cross-modal encoder to achieve enhanced image features $\boldsymbol{F}_{\text{img}}$ and text features $\boldsymbol{F}_{\text{text}}$. These enhanced cross-modal features are used to initialize a large set of query embeddings, which are individually fed into a cross-modal decoder to generate object features. Finally, each object feature is passed to the box head and the classification head to predict the bounding boxes and scores. Here, the scores are defined as the similarities between an object feature and the text tokens. To adapt to RVOS, we define the binary probability of an object being referred to by the text as the maximum score of the object over all tokens.
@@ 4 ReferDINO
We illustrate our ReferDINO in Figure 2. Given a video clip of $T$ frames and a text description, we employ GroundingDINO to derive the cross-modal image feature $\boldsymbol{F}^{t}_{img}$, text feature $\boldsymbol{F}^{t}_{text}$ and object features $\mathcal{O}^{t}$ of each individual frame. Then, we feed all the object features $\{\mathcal{O}^{t}\}^{T}_{t=1}$ into object-consistent temporal enhancer4.1), which performs inter-frame object interactions and derives temporal enhanced object features $\{\tilde{\mathcal{O}}^{t}\}^{T}_{t=1}$. These features are passed through grounding-guided deformable mask decoder4.2) to produce instance masks, conditioned on the location predictions, cross-modal text features and high-resolution feature maps. To improve the efficiency without compromising performance, we introduce a confidence-aware query pruning strategy (§4.3) in the cross-modal decoder. We detail the training and inference in §4.4.
@@ 4.1 Object-consistent Temporal Enhancer
Although GroundingDINO can detect referring objects from single images, this is not reliable enough for RVOS. First, it lacks the capability to capture temporal dynamics in videos, making it unable to handle motion-related descriptions, e.g., “the cat swing its tail”. Second, video frames often contain camera motion blur and constrained perspectives, which greatly undermine the temporal consistency of GroundingDINO. To overcome these limitations, we propose object-consistent temporal enhancer upon the GroundingDINO. This module performs inter-frame object interactions under the text guidance, enabling our ReferDINO to capture effective temporal associations between visual objects and text descriptions.
As shown in Figure 3, the proposed object-consistent temporal enhancer consist of a memory-augmented tracker and a cross-modal temporal decoder. It receives two inputs: all the object embeddings $\{{\mathcal{O}}^{t}\}_{t=1}^{T}$, and all the cross-modal sentence features $\{\boldsymbol{f}_{\text{cls}}^{t}\}_{t=1}^{T}$, where $\boldsymbol{f}_{\text{cls}}^{t}$ corresponds to the [CLS] token in the text feature $\boldsymbol{F}^{t}_{\text{text}}$.
@@
Memory-augmented Tracker. Before temporal interaction, we need to align the objects across different frames with a tracker module. Different from the trackers used in previous works [5, 8] that only consider the alignment between adjacent frames, we integrate a memory mechanism for stable long-term tracking. Let $\mathcal{M}^{t}$ indicates the memory for $t$-th frame, and $\mathcal{M}^{1}={\mathcal{O}}^{1}$. Our memory-augmented tracker consist of two steps: object alignment and memory updating. In the first step, we compute the cosine similarity between $\mathcal{M}^{t-1}$ and ${\mathcal{O}}^{t}$ as the assignment cost, and apply the Hungarian algorithm [17] to align the objects with the memory:
@@
where ${\hat{\mathcal{O}}}^{t}$ indicates the aligned object embeddings. In the second step, these embeddings are used to update the memory in a momentum-based manner. Meanwhile, we incorporate the text relevance to adaptively prevent object-invisible frames from disturbing the long-term memory. Formally, the memory is updated as follows:
@@
where $\alpha$ is the momentum coefficient, and $\boldsymbol{c}\in\mathbb{R}^{N_{s}}$ is the cosine similarity between ${\hat{\mathcal{O}}}^{t}\in\mathbb{R}^{N_{s}\times d}$ and its sentence embedding $\boldsymbol{f}_{\text{cls}}^{t}\in\mathbb{R}^{d}$.
Cross-modal Temporal Dehancer. This module takes the frame-wise sentence embeddings as frame proxies to perform inter-frame interaction and video-level object enhancement. Specifically, this module comprises $L_{t}$ blocks. In each block, given the aligned object embeddings $\{\hat{\mathcal{O}}^{t}\}_{t=1}^{T}$ and the sentence embeddings $\{\boldsymbol{f}^{t}_{\text{cls}}\}_{t=1}^{T}$, we employ self-attention along the temporal dimension to achieve inter-frame interaction. Next, we extract video-level object information with a cross-attention module, which takes the sentence embeddings as query and the object embeddings as key and value, deriving video-level object representations $\{\mathcal{O}_{v}^{t}\}_{t=1}^{T}$. These representations, containing effective temporal information, are used to enhance the frame-wise object embeddings as follows:
@@
4.2 Grounding-guided Deformable Mask Decoder
To fully leverage GroundingDINO’s pretrained capabilities for object segmentation, we customize a novel grounding-guided deformable mask decoder. This decoder integrates the pretrained representations and grounding predictions from GroundingDINO to iteratively refine the mask embeddings, enabling accurate and consistent object segmentation throughout the video.
For simplicity, we use $\tilde{\boldsymbol{o}}$ to represent an arbitrary object in $\tilde{\mathcal{O}}^{t}$. Before the mask decoder, we feed $\tilde{\boldsymbol{o}}$ into the bounding box to predict its bounding box $\boldsymbol{b}\in\mathbb{R}^{4}$, where $\boldsymbol{b}=\{b_{x},b_{y},b_{w},b_{h}\}$ encodes the normalized box center coordinates, box height and width. Meanwhile, we employ a feature pyramid network (FPN) upon the cross-modal image features $\boldsymbol{F}_{\text{img}}$, producing a high-resolution feature map $\boldsymbol{F}_{\text{seg}}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times d}$, where $H$ and $W$ denote the height and width of the raw video frames.
Then, we input these features and box predictions into the grounding-guided deformable mask decoder for object refinement and mask generation. This mask decoder consists of $L_{m}$ blocks, with each block comprising two components: deformable cross-attention [39] and vanilla cross-attention. The deformable cross-attention integrates the pretrained grounding knowledge by taking $\tilde{\boldsymbol{o}}$ as query, $\boldsymbol{F}_{\text{seg}}$ as memory and the box center $\{b_{x},b_{y}\}$ as reference point. This mechanism efficiently aggregates the spatial information around predicted locations for object refinement. The vanilla cross-attention is used to integrate the text conditions, by taking $\tilde{\boldsymbol{o}}$ as query and $\boldsymbol{F}_{\text{text}}$ as key and value. These two components work together to ensure that object segmentation is tightly coupled with both the text prompts and visual grounding. Finally, for each object query, we obtain a refined mask embedding $\boldsymbol{o}_{m}\in\mathbb{R}^{d}$, which is then dot-producted with the high-resolution feature map $\boldsymbol{F}_{\text{seg}}$ to generate an instance mask $\boldsymbol{m}$.

### v2

@@ Abstract
Referring video object segmentation (RVOS) aims to segment target objects throughout a video based on a text description. This is challenging as it involves deep vision-language understanding, pixel-level dense prediction and spatiotemporal reasoning. Despite notable progress in recent years, existing methods still exhibit a noticeable gap when considering all these aspects. In this work, we propose ReferDINO, a strong RVOS model that inherits region-level vision-language alignment from foundational visual grounding models, and is further endowed with pixel-level dense perception and cross-modal spatiotemporal reasoning. In detail, ReferDINO integrates two key components: 1) a grounding-guided deformable mask decoder that utilizes location prediction to progressively guide mask prediction through differentiable deformation mechanisms; 2) an object-consistent temporal enhancer that injects pretrained time-varying text features into inter-frame interaction to capture object-aware dynamic changes. Moreover, a confidence-aware query pruning strategy is designed to accelerate object decoding without compromising model performance. Extensive experimental results on five benchmarks demonstrate that our ReferDINO significantly outperforms previous methods (e.g., +3.9% $\mathcal{J}\&\mathcal{F}$ on Ref-YouTube-VOS) with real-time inference speed (51 FPS).
@@ 1 Introduction
Figure 1: (a) The SOTA method [23] in RVOS struggles to distinguish similar objects properly based on compositional attribute descriptions (e.g., shapes + colors), while our ReferDINO overcomes this limitation. (b) GroundingDINO [22] fails to identify the target cat based on the motion reference “swinging its tail”, while our ReferDINO corrects the prediction through cross-modal spatiotemporal reasoning and generates precise object masks.
@@
Despite significant progress in recent years, existing RVOS models [3, 40, 23, 25, 45, 9] still suffer from many common problems. For example, they often struggle to handle the queries involving complex appearance, relative locations and attributes, as shown in Figure 1 (a). These problems are mainly caused by the insufficient vision-language capabilities of current models, which, in turn, result from the limited scale and diversity of available RVOS data, as densely annotating video-text data is extremely expensive.
The recent progress in visual grounding field [22, 19, 5] offers unique potential to address this data limitation. Benefiting from large-scale pretraining on image-text data, the representative foundational models (e.g., GroundingDINO [22]) showcase profound capabilities in object-level vision-language understanding. However, simply employing them to tackle RVOS task is impracticable, as shown in Figure 1 (b). There are mainly two challenges: (1) These models well-designed for region-level regression lack the ability of pixel-level dense prediction. (2) Although they excel in understanding static attributes that can be observed in single frames (e.g., “the white cat”), they fail to identify the target object described by dynamic attributes (e.g., “the cat swinging its tail”).
To address these challenges, we propose ReferDINO, a strong end-to-end RVOS approach that inherits static object perception from GroundingDINO [22], and is further endowed with pixel-wise dense segmentation and spatio-temporal reasoning capabilities. First, we elaborate a grounding-guided deformable mask decoder to produce high-quality object masks on each frame. Instead of simply adding a mask prediction branch in parallel with the original box prediction branch [18, 40, 23, 9], our mask decoder cascades the two branches as a grounding-deformation-segmentation pineline. It utilizes the pretrained box prediction as location prior, progressively refining the mask prediction through deformable attention mechanisms. This process is differentiable, enabling mask learning to feed back into the box prediction branch for collaborative task learning. Second, we present an object-consistent temporal enhancer that injects pretrained time-varying text features into inter-frame interaction for cross-modal temporal reasoning. These two modules work together to bridge the gap between visual grounding and RVOS, effectively overcome the limitations of both foundational visual grounding models and existing RVOS methods, as shown in Figure 1.
Another limitation of these foundational models is their huge computational overhead, which is commonly unacceptable for video-scale training and inference. To address this, we further propose a confidence-aware query pruning strategy to reduce per-frame computations without compromising pretrained knowledge, by progressively identifying and pruning the low-confidence object queries.
Extensive experiments on five public RVOS benchmarks demonstrate that ReferDINO significantly outperforms state-of-the-art (SOTA) methods. Moreover, compared to other baselines built upon GroundingDINO, our approach achieves remarkable performance improvements, with +2.6 $\sim$ 3.6% $\mathcal{J}\&\mathcal{F}$ on the Ref-Youtube-VOS dataset and and +1.4 $\sim$ 4.6% $\mathcal{J}\&\mathcal{F}$ on the MeViS. In addition, our query pruning strategy reduces 40.6% FLOPs and 41.3% memory usage, achieving comparable performance with a 10$\times$ speedup (4.9 $\rightarrow$ 51.0 FPS). These results demonstrate ReferDINO’s potential for real-time video applications.
Overall, our contributions are summarized as follows:
  • We propose ReferDINO, a strong RVOS approach that inherits static object perception from pretrained foundational models and extends their capabilities to pixel-wise dense segmentation and spatiotemporal reasoning through a grounding-guided deformable mask decoder and an object-consistent temporal enhancer.
  • We introduce a confidence-aware query pruning strategy, which significantly improves training and inference efficiency without compromising performance.
  • ReferDINO establishes new SOTA performance on five benchmarks, and our proposed components provide significant improvements in performance and efficiency.
@@ 2 Related Works
Figure 2: Overall architecture of ReferDINO. Modules colored in blue are borrowed from GroundingDINO, while those in red are newly introduced in this work. First, a confidence-aware query pruning strategy is employed to progressively prune low-confidence queries, deriving only a compact set of important object features $\mathcal{O}^{t}$. After collecting the object features from all frames $\{{\mathcal{O}}^{t}\}_{t=1}^{T}$, the object-consistent temporal enhancer performs cross-modal temporal reasoning by injecting time-varying text embeddings. Finally, our grounding-guided deformable mask decoder takes the box predictions as location priors to refine the mask features through deformation attention and cross-modal attention mechanisms. Note that this process is end-to-end differentiable, enabling segmentation gradients to propagate backward into the box head for collaborative learning. Best viewed in color.
@@
Referring Video Object Segmentation. RVOS [7, 6, 31] aims to segment objects throughout the video based on text descriptions. MTTR [3] firstly introduces the DETR paradigm [4] into RVOS. Furthermore, ReferFormer [40] proposes to produce the queries from the text description. On the top of this pipeline, follow-up works [8, 25, 23, 45] focus on modular improvements. Despite notable progress on specific datasets, these models are limited by insufficient vision-language understanding, and often struggle in unseen objects or scenarios. Recently, some works [1, 13] attempt to utilize GroundingDINO to identify objects in single frames and then apply SAM2 [28] to generate dense masks. However, such a manner of model ensemble is inefficient and non-differentiable, preventing further refinement of models’ task-specific capabilities. In contrast, our ReferDINO is an end-to-end adaptation approach that benefits from both open-world knowledge of GroundingDINO and specific knowledge from RVOS data.
Visual Grounding Foundational Model. Visual grounding [10, 12, 20, 30, 35, 36, 21, 34] aims to localize object regions in an image for a given text. Recent works [19, 5, 44, 22] focus on unifying detection and image-text datasets to increase the training vocabulary at scale. For example, GroundingDINO [22] incorporates grounded pretraining into DINO [46] to achieve impressive performance. These models showcase strong object-level vision-language understanding, and have been explored in many downstream works [38, 27, 43, 47, 29]. However, in most of these works, foundational models are primarily used as preprocessing tools to extract object regions, and their further potential in end-to-end task adaptation remains largely under explored. Recently, Video-GroundingDINO [38] extend GroundingDINO to produce temporal boundaries by simply inserting several temporal self-attention modules. To the best of our knowledge, our work is the first end-to-end approach that adapts GroundingDINO to RVOS task. Besides, our designs upon foundational models can provide insights for other downstream fields.
@@ 3 Background: GroundingDINO
Our approach builds on a visual grounding foundational model, GroundingDINO [22], which we briefly revisit here. GroundingDINO is a DETR-based object detector, which introduces language into an object detector to achieve visual grounding. It mainly consists of an image backbone, a text backbone, a cross-modal encoder-decoder Transformer architecture, a box head, and a classification head.
Given an image-text pair, GroundingDINO adopts the dual backbones to extract vanilla features, which are then fed into the cross-modal encoder to derive enhanced image features $\boldsymbol{F}_{\text{img}}$ and text features $\boldsymbol{F}_{\text{text}}$. These enhanced cross-modal features are used to initialize a large set of query embeddings $\mathcal{Q}$, which are individually fed into a cross-modal decoder to generate object features $\mathcal{O}$. Note that the shapes of $\mathcal{O}$ and $\mathcal{Q}$ are equal. Finally, each object feature is passed to the box head and the classification head to predict bounding boxes and scores. Here, the scores are defined as the similarities between each object feature and the text tokens. For RVOS, we define the binary classification probability as the maximum score over all tokens.
@@ 4 ReferDINO
We illustrate our ReferDINO in Figure 2. Apart from GroundingDINO, our model consists of three main components: grounding-guided deformable mask decoder (§4.1), object-consistent temporal enhancer (§4.2) and confidence-aware query pruning (§4.3). First, given a $T$-frame video and a language query, GroundingDINO is separately applied to each frame to obtain the object features. To improve per-frame efficiency, we use the confidence-aware query pruning strategy to progressively remove low-confidence object queries, achieving a very compact set of object features. Then, collecting the object features of all frames, we utilize object-consistent temporal enhancer to perform cross-modal temporal reasoning and ensure temporal consistency. After that, we employ the grounding-guided deformable mask decoder, which takes each object’s box prediction as location condition, progressively refining its feature through the deformable cross-attention and cross-modal attention. Finally, the output feature corresponding to each object is used to produce a mask sequence $\{\mathbf{m}^{t}\}^{T}_{t=1}$ by dot-producting with the per-frame feature maps.
4.1 Grounding-guided Deformable Mask Decoder
This decoder takes an object feature as input and generates corresponding pixel-wise predictions on each frame. Instead of simply adding a mask head in parallel with the original box head, our mask decoder formulates the box and mask predictions as a grounding-deformation-segmentation pineline. It utilizes the box prediction as location prior to iteratively enhance the mask predictions, and incorporates the text prompts for further refinement. Moreover, this process is differentiable, enabling mask learning to feed back into the box head for collaborative task learning.
For simplicity, we use $\tilde{\boldsymbol{o}}$ to represent an arbitrary object feature. Before the mask decoder, we feed $\tilde{\boldsymbol{o}}$ into the box head to predict its bounding box $\boldsymbol{b}\in\mathbb{R}^{4}$, where $\boldsymbol{b}=\{b_{x},b_{y},b_{w},b_{h}\}$ encodes the normalized box center coordinates, box height and width. Meanwhile, we employ a feature pyramid network (FPN) upon $\boldsymbol{F}_{\text{img}}$, producing a high-resolution feature map $\boldsymbol{F}_{\text{seg}}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times d}$, where $H$ and $W$ denote the height and width of the raw video frames.
Then, we input these features and box predictions into the mask decoder for location-aware object refinement and mask generation. Our mask decoder consists of $L_{m}$ blocks, with each block comprising two components: deformable cross-attention [48] and cross-modal attention. When applying attention on image feature maps (termed memory), deformable cross-attention aggregates the features by adaptively sampling a small set of sampling points around a reference point. Unlike the vanilla deformable cross-attention [48] that generates reference points through MLPs, we directly use the predicted box center $\{b_{x},b_{y}\}$ as the reference point. Specifically, we treat $\tilde{\boldsymbol{o}}$ as the query, $\boldsymbol{F}_{\text{seg}}$ as the memory, and the normalized box center $\{b_{x},b_{y}\}$ as the reference point. The sampling process is implemented by bilinear interpolation [16], thus it is end-to-end differentiable. In this way, the pretrained grounding knowledge is naturally incorporated to refine the mask prediction, and the segmentation gradients are also backward to optimize the object grounding. A concern is that noisy features (e.g. background) can also be involved during adaptive sampling, which potentially compromises the mask quality. To mitigate this, we further apply cross-modal attention to integrate the text conditions, by taking $\tilde{\boldsymbol{o}}$ as query and $\boldsymbol{F}_{\text{text}}$ as key & value. These two attention mechanisms work together to ensure that mask prediction is tightly coupled with both text prompts and object locations. Finally, for each object query, we obtain a refined mask embedding $\boldsymbol{o}_{m}\in\mathbb{R}^{d}$, which is then dot-producted with the high-resolution feature map $\boldsymbol{F}_{\text{seg}}$ to generate an instance mask $\boldsymbol{m}$.
Discussion. The dynamic mask head [33] also integrates object location information into mask prediction by separately concatenating the feature map with each object’s relative coordinates, and it has been adopted in many previous works [3, 40, 23, 39, 25, 45]. However, the memory cost of this approach is dramatically high, due to the storage demand of per-object high-resolution feature maps. This is particularly unacceptable for foundational models, which typically uses a large number of object queries. In contrast, our mask decoder efficiently integrates the location information of different objects by location-guided sampling on a shared feature map. This design requires no extra memory burden, making our mask decoder particularly suitable for adaptation of foundational models.
@@ 4.2 Object-consistent Temporal Enhancer
Although GroundingDINO can detect referring objects from single images, this is not reliable enough for RVOS. First, it fails to understand the dynamic attributes in descriptions (e.g., action or temporal relationship). Second, video frames often contain camera motion blur and constrained perspectives, greatly undermining its temporal consistency. Therefore, we introduce an object-consistent temporal enhancer to enable corss-modal temporal reasoning. As shown in Figure 3, this module consists of a memory-augmented tracker and a cross-modal temporal decoder. It receives two inputs: all the object embeddings $\{{\mathcal{O}}^{t}\}_{t=1}^{T}$, and time-varying sentence features $\{\boldsymbol{f}_{\text{cls}}^{t}\}_{t=1}^{T}$, where $\boldsymbol{f}_{\text{cls}}^{t}$ corresponds to the [CLS] token in $\boldsymbol{F}^{t}_{\text{text}}$.
@@
Memory-augmented Tracker. Before temporal interaction, we align the objects across different frames with a tracker mechanism. Let $\mathcal{M}^{t}$ indicates the memory for $t$-th frame, and $\mathcal{M}^{1}={\mathcal{O}}^{1}$. Our tracker consists of two steps: object alignment and memory updating. In the first step, we compute the cosine similarity between $\mathcal{M}^{t-1}$ and ${\mathcal{O}}^{t}$ as the assignment cost, and apply the Hungarian algorithm [17] to align the objects with the memory:
@@
where ${\hat{\mathcal{O}}}^{t}$ indicates the aligned object embeddings. In the second step, these embeddings are used to update the memory in a momentum manner. Meanwhile, we incorporate the text relevance to prevent object-invisible frames from disturbing the long-term memory. Formally, the memory is updated as follows:
@@
where $\alpha$ is the momentum coefficient, and $\boldsymbol{c}\in\mathbb{R}^{N_{s}}$ is the cosine similarity between ${\hat{\mathcal{O}}}^{t}\in\mathbb{R}^{N_{s}\times d}$ and $\boldsymbol{f}_{\text{cls}}^{t}\in\mathbb{R}^{d}$.
Cross-modal Temporal Decoder. This module takes the time-varying text embeddings as frame proxies to perform inter-frame interaction and temporal enhancement. Specifically, this module comprises $L_{t}$ blocks. In each block, given the aligned object embeddings $\{\hat{\mathcal{O}}^{t}\}_{t=1}^{T}$ and the sentence embeddings $\{\boldsymbol{f}^{t}_{\text{cls}}\}_{t=1}^{T}$, we employ self-attention along the temporal dimension to achieve inter-frame interaction. Next, we extract dynamic information with a cross-attention module, which takes the sentence embeddings as query and the object embeddings as key and value, deriving temporal-enhanced object features $\{\mathcal{O}_{v}^{t}\}_{t=1}^{T}$. These features containing effective temporal information are used to enhance the frame-wise object embeddings as follows:
@@
Discussion. Our temporal enhancer improves existing temporal modules of RVOS models in two aspects. 1) Previous works either ignore object tracking [37, 23, 25, 32] or only consider it in adjacent frames [6, 39, 9], while we employ a memory-augment tracker for stable long-term consistency. 2) In cross-modal temporal interaction, prior methods [40, 23, 32, 9] always use a static text embedding from the text encoder (e.g., BERT) for all objects across different frames, while our temporal enhancer uses time-varying text embeddings to better capture fine-grained temporal dynamics. These two designs make our temporal enhancer fully exploit the pretrianed object knowledge and cross-modal features from foundational models for effective spatiotemporal reasoning and object consistency.